OpenStudy (kkutie7):

I believe this series is convergent but i'm not sure what test to use to figure it out.

OpenStudy (kkutie7):

\[\sum_{2}^{\infty}\frac{(-1)^{n}}{(\ln(n))^{2}}\]

OpenStudy (anonymous):

alternates, right?

OpenStudy (ac3):

alternating would probably work

OpenStudy (kkutie7):

Yes I think so

OpenStudy (kkutie7):

alright

OpenStudy (anonymous):

since it alternates \((-1)^n\) tells you that, all that is necessary is that the absolute value of the terms are a) decreasing and b) go to zero

OpenStudy (anonymous):

which they do

OpenStudy (kkutie7):

thank you

OpenStudy (anonymous):

yw